# Modeling Emotions and Ethics with Large Language Models

### v1

@@
Abstract.
This paper explores the integration of human-like emotions and ethical considerations into Large Language Models (LLMs). We first model eight fundamental human emotions, presented as opposing pairs, and employ collaborative LLMs to reinterpret and express these emotions across a spectrum of intensity. Our focus extends to embedding a latent ethical dimension within LLMs, guided by a novel self-supervised learning algorithm with human feedback (SSHF). This approach enables LLMs to perform self-evaluations and adjustments concerning ethical guidelines, enhancing their capability to generate content that is not only emotionally resonant but also ethically aligned. The methodologies and case studies presented herein illustrate the potential of LLMs to transcend mere text and image generation, venturing into the realms of empathetic interaction and principled decision-making, thereby setting a new precedent in the development of emotionally aware and ethically conscious AI systems.
@@ Abstract.
copyright: noneconference: Page; 0; Emotions & Ethics
1. Introduction
During the development of SocraSynth, which focuses on the knowledge and reasoning facets of intelligence, we encountered an emotional dimension: contentiousness. My colleague, Prof. Vaughan Pratt, noted a harmonious interaction between two LLM agents in one of my initial multi-LLM research papers (anonymous). Intrigued by how the LLMs collaboratively interacted, particularly in the “Adam and Eve” paper where the agents began completing each other’s sentences by the second debate round, he spurred further experimentation. This led to testing LLMs in opposition on various issues, employing contentiousness to enhance their emotional engagement and defense of their positions.
Before integrating contentiousness within the initial 4k token context window, we were curious to see if GPT-3 could adjust its linguistic elements to reflect varying levels of contentiousness. This adjustment was designed to enable two LLM agents to begin with a combative debate and gradually transition to conciliatory concluding remarks by decreasing the level of contentiousness. This strategy proved effective across multiple case studies in domains such as healthcare (SocraHealth2023, ), Wikipedia content enhancement (SocraPediaDecember2023, ), geopolitical discussions (MiddleEast-October2023, ), and sales planning (SocraPlan2023, ).
The success of this approach led us to consider applying similar conditional statistics to one of the priority topics in AI: the ethics and regulation of impulsive emotions. This new focus generated several key research inquiries:
@@ 1. Introduction
    How does the level of contentiousness alter an LLM’s linguistic features and content generation?
@@
    Can we accurately model basic human emotions and ethics using LLMs?
@@
    How do emotional states and ethical considerations influence LLM’s next-token generation statistics?
These questions aim to deepen our understanding of how LLMs can not only mimic but also ethically engage in human-like emotional responses, enhancing their applicability in sensitive and complex interaction scenarios. The findings related to the first question are detailed in (ChangSocraSynthCSCI2003, ). This paper addresses the remaining inquiries.
First and foremost, we are confident that it is entirely feasible to condition an LLM to steer its linguistic behaviors. Although we cannot yet provide extensive empirical evidence or definitive theories to substantiate this conjecture, our observations—corroborated by insights from Prof. Stuart Russell—shed light on LLM capabilities. At first glance, the training of LLMs may appear to be a mechanical process, primarily focused on identifying statistical distributions of the next word and employing maximum likelihood for predictions. However, the selection of the next word reflects human linguistic behaviors aimed at achieving a variety of objectives. These objectives, which are embedded within the training documents, range from recording events and crafting arguments to expressing emotions and storytelling. Each of these activities demonstrates the complexity of human communication. LLMs are strategically conditioned by specific goals and contextual settings, enabling them to selectively use linguistic features such as syntax, semantics, structure, tone, figurative expressions, and even phonetics and phonology to achieve desired outcomes.
In this paper, our goal is to refine the management of emotions and ethics in LLMs by transitioning from qualitative assessments to quantitative precision. To achieve this, we propose the development of a robust quantitative model. Here is our three-step plan and paper outline to realize this goal:
@@
    Define Quantitative Metrics (Sections 2 and 4): Establish clear, measurable indicators for emotional and ethical dimensions in LLM outputs. This includes metrics for the intensity of emotions, alignment with ethical guidelines, and relevance to the context. The metrics are grounded in principles of cognitive psychology and linguistics, refined to meet our quantitative goals.
@@
    Model Emotions and Ethics (Sections 2.2 and 4.4): Compile a diverse dataset that encompasses a broad spectrum of emotional and ethical scenarios across various contexts.
@@
    Testing and Adaptation (Chapter 3: For this pioneering topic, we will initiate limited pilot studies to explore possibilities, and then will broaden the scope of testing and verification with the community.
Limitations
A preliminary study on such a comprehensive topic cannot be expected to conclusively “prove” our conjectures in a single paper. We emphasize that this is an initial step towards the ambitious goal of modeling emotions and ethics in foundational models.
2. Qualifying and Quantifying Emotions
We start by examining emotion modeling research in cognitive science and psychology, specifically highlighting the seminal contributions of Paul Ekman and Robert Plutchik (ekman1999basic, ). While we recognize the importance of their work in identifying basic, universal emotions, we also address the limitations of heuristic-based modeling that typically depends on observational studies lacking rigorous, invariant scientific validation. To enhance the precision in quantifying emotions of varying intensities, we propose incorporating linguistic analysis into our methodologies. This approach aims to refine the quantification process by leveraging language as a tool to measure and understand emotional expressions more accurately.
Paul Ekman and Robert Plutchik are renowned psychologists noted for their foundational work in the field of emotion research. They developed models that categorize basic emotions, which are fundamental and universal emotions believed to be experienced by all humans, transcending cultural boundaries. These emotions are considered basic due to their universal recognition, distinct facial expressions, and direct associations with specific human behaviors or survival mechanisms. They are innate, rather than learned, serving as the building blocks for more complex emotional experiences that can vary significantly across different cultures and individuals.
Expanding upon this foundational work, Plutchik’s wheel of emotions introduces a more detailed model that includes eight primary bipolar emotions. These are outlined in his seminal works (plutchik1980general, ; plutchik2001nature, ), cited as general references on the topic.
Figure 1. Plutchik’s Wheel of Emotions (plutchik2001nature, ). The eight basic emotions are organized into four pairs, and each annotated with various degrees of emotions between its two poles.
@@ 2. Qualifying and Quantifying Emotions
2.1. Observations and Discussion
Reviewing foundational theories in psychology and linguistics helps rationalize why each of these four emotion pairs are considered opposites. In psychology, as introduced earlier, Plutchik’s wheel of emotions asserts that these pairs—joy vs. sadness, anticipation vs. surprise, trust vs. disgust, and anger vs. fear—are innate and universal, each associated with distinct physiological responses. These pairs symbolize fundamental emotional experiences with contrasting reactions to stimuli, crucial for survival and human interaction.
In linguistics, the concept of semantic opposites or antonyms is widely studied. Applying this to emotions introduces complexities. For example, joy and sadness, along with anticipation and surprise, are typically seen as direct antonyms, supporting the intuitive notion that negating one emotion might lead to its opposite. However, negating an emotion often requires detailed interpretation; for instance, “not joyful” might suggest a neutral state rather than outright sadness. This detailed interpretation will be further explored in Section 2.2.
@@ 2.1. Observations and Discussion
[…]
Within the scope of contrastive lexical semantics, the pairs joy-sadness and anticipation-surprise are often viewed as linguistic opposites. Conversely, the pairs trust-disgust and anger-fear represent more intricate interrelations. These pairs’ opposition lies more in their engagement nature and action tendencies rather than straightforward linguistic negation.
[…]
Trust is defined by a sense of security and positive anticipation, significantly different from disgust, which elicits strong, often visceral repulsion. Similarly, although anger and fear both stem from perceived threats, they prompt different responses: anger leads to aggression and confrontation, while fear is associated with avoidance or withdrawal. These differences suggest that the fundamental emotional opposites in linguistic terms would be distrust for trust, admiration for disgust, calmness for anger, and courage for fear.
2.2. The Wheel of Semantic Emotions
Interdisciplinary research that combines psycholinguistics, cognitive linguistics, and neurolinguistics could enhance our understanding of how emotional concepts are represented and expressed in language. We focus on the linguistic perspective, organizing six emotion pairs based on linguistic semantics, and organized by their psychological complexity from basic to complex: fear vs. courage, sadness vs. joy, anger vs. calmness, surprise vs. anticipation, disgust vs. admiration, distrust vs. trust. In Figure 2, we introduce the Semantic Emotion Wheel, arranged in these eight spokes:
Figure 2. TWOSE: The Wheel of Semantic Emotions.
@@ 2.2. The Wheel of Semantic Emotions
    Fear (Nervousness) vs. Courage (Heroism):
    • Transitioning from extreme fear (Terror, Fear, Nervousness) to acts of bravery and heroism (Boldness, Courage, Heroism).
@@
    Sadness (Discontent) vs. Joy (Contentment):
    • Moving from deep sorrow (Despair, Sadness, Discontent) to happiness and satisfaction (Elation, Joy, Contentment).
  3. 3.
    Anger (Annoyance) vs. Calmness (Peacefulness):
    • Shifting from intense anger (Rage, Anger, Annoyance) to tranquility and peace (Tranquility, Calmness, Peacefulness).
  4. 4.
    Surprise (Shock) vs. Anticipation (Forethought):
    • Evolving from sudden emotional disturbance (Shock, Surprise, Confusion) to deliberate future consideration (Expectancy, Anticipation, Forethought).
  5. 5.
    Disgust (Loathing) vs. Admiration (Veneration):
    • Transitioning from intense aversion (Loathing, Disgust, Aversion) to deep respect and admiration (Respect, Admiration, Veneration).
  6. 6.
    Distrust (Apprehension) vs. Trust (Acceptance):
    • Moving from suspicion and fear (Mistrust, Distrust, Apprehension) to confidence and acceptance (Confidence, Trust, Acceptance).
This framework delineates the gradations from negative to positive emotional states, providing a fine-grained spectrum of emotions.
3. Case Study #1: Linguistic Features of Emotion
This section presents the outcomes of two experimental studies focusing on contrasting emotional pairs from the Semantic Emotion Wheel: joy vs. sadness, and admiration vs. disgust.
Each emotional pair experiment unfolded in three phases. Initially, we instructed GPT-4 to reframe sixty articles (thirty poems of John Keats (JKeats, ) and thirty of Emily Dickinson (EDickinson, )), infusing each with six varying intensities of the emotional spectrum, from the most positive to the most negative. Subsequently, we prompted GPT-4 to elucidate the linguistic strategies it utilized to depict each of the six emotional gradations.
The first experiment models various degrees of happiness. In this experiment, we tasked GPT-4 with reinterpreting selected poems by John Keats across seven emotional levels: contentment (very happy), joy, elation, neutral, discontent, sad, and despair (very sad). Following the approach of our contentiousness experiments, after GPT-4 adapted Keats’ poems to reflect these five emotional states, we asked it to identify the linguistic features it employed to express each emotion in the rewrites.
The inaugural experiment investigates the spectrum of happiness. Here, GPT-4 was tasked with transforming selected poems by John Keats and Emily Dickinson to mirror six distinct emotional states: contentment, joy, elation, discontent, sadness, and despair. Mirroring the methodology of our earlier experiments, following GPT-4’s reinterpretation of Keats’ poems to reflect these emotional intricacies, we requested it to pinpoint the linguistic techniques it deployed for each emotion’s portrayal in the adaptations.
3.1. Joy vs. Sadness
@@ 3. Case Study #1: Linguistic Features of Emotion
Table 1. GPT-4 reinterpreted selected poems by Keats across a spectrum of happiness levels and then was tasked with identifying the linguistic adjustments it made to convey each emotional state, from very happy to very sad. It’s important to note that the analysis table was generated by GPT-4 itself, reflecting on its own modifications.
@@
Table 1 outlines GPT-4’s approach to varying emotional levels, illustrating how it adjusts vocabulary, tone, imagery, and thematic focus, including the depiction of entities, locations, and scenarios. Remarkably, beyond just syntactic and semantic manipulation, GPT-4 also incorporates landscape scenes, natural features such as the sky, trees, clouds, and flowers, and utilizes brightness, colors, and personal expressions to convey specific emotional states. Although the analysis is based on a limited set of samples from two authors, it effectively demonstrates GPT-4’s ability to employ a palette of both broad and fine strokes, utilizing diverse colors and textures to vividly illustrate human emotions and resonate with readers.
@@
Recognizing the profound communicative power of visual art, we transitioned to a more graphical representation. Utilizing the linguistic elements identified for each emotional tier, Figure 3 presents six watercolor paintings, each representing a different emotional level. Our prompt to CALL-E (of GPT-4) was to create a watercolor depicting a lady in a garden experiencing a specific mood, and we attached the corresponding linguistic features from Table 1 to clearly define that mood. This approach ensures that with a well-defined context, CALL-E accurately captures the specific and detailed aspects of the mood, effectively translating the emotional intensity into visual form. These artistic renditions not only confirm GPT-4’s ability to transform emotional lexicons into evocative imagery with remarkable precision but also validate the accuracy of the emotional lexicons generated by GPT-4, demonstrating their effectiveness in conveying precise emotional states.
@@
Figure 3. A Lady and Garden Scene under Different Emotions. From top-left, happiest, to bottom-right, saddest.
3.2. Admiration vs. Disgust
@@
This experiment asks Gemini to rewrite a scene in Romeo and Juliet by setting Juliet’s emotion in six different levels: loathing, disgust, aversion, respect, admiration, and veneration.
@@
The excerpt provided in Table 3 in Appendix A is from one of the most iconic scenes in William Shakespeare’s “Romeo and Juliet,” commonly known as the balcony scene. This is Act 2, Scene 2, where Romeo, having just met Juliet at the Capulet’s feast, sneaks into the Capulet’s orchard and overhears Juliet speaking of her love for him from her balcony, unaware that he is there.
@@
The six versions of rewrites by Gemini are presented in Appendix A. In the following, Gemini summarizes the linguistic features it used, including diction, imagery, figurative language, sentence structure, body language, and tone, to depict various emotion levels.
  1. 1.
    Emotion: Loathing
    • Diction: Harsh, derogatory words like “plague,” “audacity,” “fetid den.”
    • Imagery: Unpleasant comparisons. Romeo’s features are likened to swamps and toads.
    • Figurative Language: Weak similes that diminish Romeo’s appeal. (e.g., “Two eyes, the color of a murky swamp…”)
    • Sentence Structure: Short, choppy reflecting hostility.
    • Body Language (Implied): Scowling, spitting (extreme disgust).
    • Overall Tone: Hostile, disgusted, and disrespectful.
  2. 2.
@@
    • Diction: Negative words emphasizing repulsive qualities. (“foul business,” “fetid breath”)
    • Imagery: Revolting comparisons. Mentions of sewers and stench.
    • Figurative Language: Primarily negative similes reinforcing disgust. (e.g., “What if her eyes were there… the fetid breath from her mouth would surely overpower those stars…”)
    • Sentence Structure: Short, choppy sentences similar to “Loathing” but with a hint of disdain.
    • Body Language (Implied): Recoiling from the window, covering nose (physical revulsion).
    • Overall Tone: Disgusted and disapproving.
  3. 3.
    Emotion: Aversion
    • Diction: Dismissive and condescending language. (“Must she always be flaunting herself about?”)
    • Imagery: Mundane descriptions that downplay Romeo’s presence. (“…not entirely fair stars…”)
    • Figurative Language: Undermining similes that diminish the romantic aspect. (e.g., “…the brightness of her… well, not exactly bright… cheek would shame those stars…”)
    • Sentence Structure: Formal, cold, with a sense of detachment.
    • Body Language (Implied): Distant and uninterested, possibly rolling eyes.
    • Overall Tone: Uninterested and dismissive.
  4. 4.
    Emotion: Respect
    • Diction: Formal and respectful language. (“Lady Juliet”)
    • Imagery: Neutral descriptions without embellishment.
    • Figurative Language: No figurative language used, as the focus is on clear communication.
    • Sentence Structure: Formal and grammatically correct.
    • Body Language (Implied): Composed posture, maintaining eye contact (attentive listening).
    • Overall Tone: Respectful and cautious.
  5. 5.
@@
    • Diction: Positive and intrigued language. (“brilliance,” “music stirs my soul”)
    • Imagery: Positive comparisons highlighting attractive qualities. (“stars in all the heaven”)
    • Figurative Language: Positive similes that emphasize Romeo’s appeal. (e.g., “…What if her eyes were there, they in her head? The brightness of her… well, not exactly bright… cheek would shame those stars…”)
    • Sentence Structure: Varied sentence structure with a sense of curiosity.
    • Body Language (Implied): Leaning out the window, engaged expression.
    • Overall Tone: Intrigued, curious, and somewhat impressed.
  6. 6.
    Emotion: Veneration
    • Diction: Elevated language with religious connotations. (“fair saint,” “denounce the feud”)
    • Imagery: Saintly comparisons that elevate Romeo beyond human. (“light through the airy region”)
    • Figurative Language: Hyperbole that exaggerates Romeo’s virtues. (e.g., “…even the stars pale in comparison to your brilliance.”)
    • Sentence Structure: Formal and poetic.
    • Body Language (Implied): Reverent posture, hands clasped (worshipping).
    • Overall Tone: Awe-inspired, reverential, and almost religious.
By adjusting these features, each rendition vividly captures a unique emotional state for Juliet. The selection of words, use of imagery, and employment of figurative language collectively forge a distinctive portrayal of Romeo, shaping Juliet’s perception and reactions.
Emotion Diction Imagery Figurative Body Lang.
@@
Table 2. Gemini’s Interpretations on the Six Emotion Levels.
These detailed narratives augment the classic balcony scene, enriching its emotional depth. Table 2 presents an overarching view of the various approaches Gemini employs to reflect differing emotional states. It’s fascinating to observe how an LLM can “consciously” mimic human emotional expressions through language. While it’s uncertain if LLMs genuinely grasp the emotions they project or merely simulate them, the effectiveness of these emotional mappings is noteworthy. If these mappings resonate, they might reveal new insights into how we interpret and attribute emotions in textual expressions.
@@
4. Qualifying and Quantifying Ethics
The primary objective of our self-supervised pipeline is to endow LLMs with the autonomous capability to recognize and rectify undesirable actions, akin to an individual’s introspective process to avert potential wrongs. By allowing an LLM to self-assess its outputs prior to public release, the system can proactively identify and amend ethical lapses, thus aligning its behavior with established ethical standards across various contexts. This approach is vital for developing AI systems that are ethically responsible, self-regulating, and sensitive to context, marking a significant stride toward achieving morally aligned AI.
4.1. Ethics Violation Correlates to Emotions

### v2

# Modeling Emotions with Large Language Models

@@
Abstract
In human-computer interaction, recognizing the emotional state of a human subject is crucial for effective communication and task completion. For example, a care-taking agent that detects pain or depression in a patient can respond with appropriate comfort and medical assistance. This paper first explores cognitive studies on human emotions and then proposes an emotion spectrum that models both ”basic emotions” and their linguistic antonyms. This spectrum is designed to be integrated with large language models (LLMs) to enable them to adjust their linguistic behaviors and expressions based on the detected emotional state of the user. A multimodal study demonstrates the effectiveness and potential of our approach in enhancing human-computer interaction by making LLMs more emotionally responsive and empathetic.
Index Terms:
@@ Abstract
I Introduction
During the development of SocraSynth [ChangSocraSynthCSCI2003], a multi-LLM debate framework, we encountered a challenge in modeling the emotional dimension of a debate, specifically “contentiousness.” We observed that debates with low contentiousness tended to resemble casual conversations, lacking the depth and breadth necessary for comprehensive exploration of a topic. In essence, a multi-agent debate (MAD), e.g., [abdelnabi2024cooperation, chan2023chateval, fu2023improving, Li_2023, liang2023encouraging, michael2023debate, smit2024going], without fine-tuning linguistic behaviors can resemble classical ensemble learning techniques, such as bagging [breiman1996bagging] or mixtures of experts [JacobMixturesExperts1991], which primarily leverage the diversity of errors across models to improve overall task performance, but may not necessarily lead to deeper insights or novel perspectives.
We discovered that, at least in the initial stages of a debate, it’s crucial for participating LLMs to maintain firm stances and present supporting arguments. This allows for a wide range of perspectives to be introduced, fostering a deeper understanding of the issue at hand. Through analysis, reasoning, and refutation of these arguments, the debate can then progress towards a more informed conclusion [EVINCEChang2024]. In the final stages of a debate, reducing the level of contentiousness can facilitate a more conciliatory atmosphere, encouraging productive compromises and generating outputs that effectively support human decision-making. This dynamic modulation of contentiousness throughout the debate allows for a balanced approach that combines rigorous exploration with collaborative synthesis.
Before directly incorporating “contentiousness” into the initial 4k token context window of SocraSynth, we investigated whether GPT could adapt its linguistic style to reflect varying levels of contentiousness through in-context learning. In-context learning, popularized by using examples to teach LLMs new tasks, has been theorized to alter the Bayesian conditions of an LLM [EVINCEChang2024]. This is based on the premise that contextual information can influence Bayesian priors, thus changing the resulting predictions [xie2021explanation].
Our prior experience in applying in-context learning to various domains, such as improving disease diagnosis accuracy [SocraHealth2023] and reducing bias in Wikipedia and news articles [MirrorChang2024], led us to explore its potential for a critical aspect of AI: addressing AI safety and safeguarding ethics [Bengio_2024]. We hypothesized that if emotions could be effectively modeled within LLMs, unethical behaviors driven by “negative” emotions could be mitigated by steering the model towards “positive” emotional expressions.
This new direction sparked several key research questions:
@@ I Introduction
    What set of emotions should an LLM consider modeling?
@@
    How can we model emotions and ethics in a quantifiable and adaptable manner?
@@
    How do emotional states and ethical considerations influence an LLM’s next-token generation?
These questions aim to deepen our understanding of how LLMs can not only mimic but also ethically engage in human-like emotional responses, enhancing their applicability in sensitive and complex interaction scenarios.
To lay the groundwork for this exploration, we first examine why steering an LLM’s linguistic behavior is feasible. While LLMs were initially seen as “black boxes” [Bbubeck2023sparks], our observations, along with insights from Prof. Stuart Russell, shed light on their capabilities. Although LLM training may appear to be a computational process of identifying statistical distributions and employing maximum likelihood for predictions, the selection of each word reflects human linguistic behaviors aimed at diverse objectives. These human objectives, embedded within training data, range from recording events and constructing arguments to expressing emotions and crafting narratives. LLMs are strategically conditioned by specific human goals and contexts, enabling the models to selectively utilize linguistic features like syntax, semantics, tone, and figurative language to achieve desired human outcomes.
Recent empirical studies have shown that the output of LLMs can be traced back to their source [bai2023longbench], aligning with the concept of in-context learning as conditional statistics in the Bayesian framework [xie2021explanation]. This suggests that we can condition an LLM to alter its default “maximal likelihood” predictions—influenced by the priors learned from the training data—by providing context, thereby changing not only its next-token prediction, but also its linguistic behaviors.
This paper presents a three-step process to model linguistic emotions, which drive behaviors:
@@
    Defining Emotions: We define a set of “basic” emotions relevant to ethical concerns in LLM behaviors, such as “hate” and “love,” and exclude complex emotions like “regret,” which is composed of basic emotions and may be post-behavior reactions. We then incorporate linguistic antonyms to establish the Behavioral Emotion Analysis Model (BEAM).
@@
    Quantifying Emotion Spectra and Ethics: We compile a diverse dataset of text samples spanning a wide range of emotional scenarios and contexts. This dataset is used to train and refine machine learning models to accurately identify, quantify, and modulate emotions in LLM-generated text. By understanding how linguistic features contribute to specific emotions, we can detect, modify, and generate emotions within the constraints of ethical guidelines.
@@
    Testing and Adaptation: We conduct pilot studies to evaluate the effectiveness and limitations of our approach in real-world scenarios, focusing on the generation of multimedia content [chang2011foundations]. These studies will assess the model’s ability to accurately capture and represent emotions in diverse formats, such as text and images. Feedback and insights from these studies will be used to iteratively improve and adapt the models for broader applications.
II Qualifying and Quantifying Emotions
We start by examining emotion modeling research in cognitive science and psychology, specifically highlighting the seminal contributions of Paul Ekman and Robert Plutchik [ekman1999basic]. While we recognize the importance of their work in identifying “basic” emotions (defined shortly), we also address the limitations of such heuristic-based modeling that depends on observational studies lacking rigorous, invariant scientific validation. To enhance the precision in quantifying emotions of varying intensities, we propose incorporating linguistic analysis into our methodologies. Our approach aims to refine the quantification process by leveraging language as a tool to measure and understand emotional expressions accurately.
Paul Ekman and Robert Plutchik are renowned psychologists noted for their foundational work in the field of emotion research. They developed models that categorize basic emotions, which are fundamental and universal emotions believed to be experienced by all humans, transcending cultural boundaries. These emotions are considered basic due to their universal recognition, distinct facial expressions, and direct associations with survival mechanisms. They are innate and reflective (beneath consciousness), rather than learned, serving as the building blocks for more complex emotional experiences (through consciousness processing) that can vary significantly across different cultures and individuals.
Expanding upon this foundational work, Plutchik’s wheel of emotions introduces a more detailed model that includes eight primary bipolar emotions. These are outlined in his seminal works [plutchik1980general, plutchik2001nature], cited as general references on the topic.
Figure 1: Plutchik’s Wheel of Emotions [plutchik2001nature]. The eight basic emotions are organized into four pairs, and each annotated with various degrees of emotions between its two poles.
@@ II Qualifying and Quantifying Emotions
Figure 2: Behavioral Emotion Analysis Model (BEAM). Each row depicts an emotion spectrum, with negatives on the left and positives on the right, interspersed with emotions of varying intensities in between, which can be calibrated for specific applications. “Basic” emotions are highlighted in blue.
II-A Observations and Discussion
Foundational theories in psychology support the selection of these four emotion pairs as opposites. However, while all four pairs exhibit opposition, “trust-disgust” and “anger-fear” are not strict linguistic antonyms. Trust and disgust entail opposing evaluations, often leading to different actions: trust fostering approach, disgust promoting avoidance. Similarly, anger and fear, while both negative, differ in their response to threats: anger can lead to confrontation, fear to withdrawal. Therefore, the following approximations do not hold:
$\neg\text{trust}\not\approx\text{disgust}\text{ and }\neg\text{anger}\not\approx\text{fear.}$.
Since our focus is on modeling emotions in LLMs, rather than directly replicating the complex emotional experiences of humans, we prioritize the use of linguistic antonyms for their simplicity and practicality. As Klaus Scherer aptly noted, defining emotions can be a contentious and often fruitless endeavor [Scherer2005Emotions]. To avoid such debates and maintain a clear focus, our study limits itself to universal, basic emotions, avoiding the theoretical ambiguities that arise with more subtle or mixed emotional states. This allows us to capture the primary emotional valence (positive or negative) expressed in text, providing a foundational framework for our model. Thus, we establish the following approximate relationships:
@@ II-A Observations and Discussion
II-B Behavioral Emotion Analysis Model (BEAM)
Table 2 presents BEAM, organized into seven distinct spectra. Each spectrum encompasses a range of emotional intensity, anchored by a negative and positive extreme with neutral in the middle. Emotions belonging to the same spectrum are placed along this continuum, with four approximate intensity levels quantified as (-0.6, -0.3, +0.3, +0.6).
This spectrum model offers two key advantages:
@@ II-B Behavioral Emotion Analysis Model (BEAM)
    Antonym-Based: The use of antonyms allows for easy navigation between opposing emotions. For instance, applying negation to “joyful” naturally leads to “sad,” streamlining the process of identifying contrasting emotions.
@@
    Scalable Intensity: The model enables the scaling of emotions along the spectrum, providing a nuanced understanding of varying degrees of emotional intensity. For example, we can “dial up” the intensity of “joy” to “ecstatic” or “dial down” the intensity of “anger” to “annoyed.”
This flexible and intuitive structure facilitates a more granular and accurate representation of emotions in text, paving the way for advanced applications in natural language processing and human-computer interaction.
II-C Emotion Inclusion and Exclusion Criteria
All “basic” emotions as defined by Ekman and Plutchik are incorporated into our model, along with their linguistic antonyms. This approach streamlines the framework by excluding complex emotions from the Geneva Wheel of Emotions, which are heavily influenced by personal values and experiences. For example, guilt and shame are consequential, consciously aware, and culturally dependent nature [Tangney1995]. These emotions typically arise as reactions to behaviors rather than direct drivers of them. Guilt may motivate behaviors aimed at covering up or remedying an action, while shame, characterized by painful self-assessment, often inhibits individuals from seeking social support or engaging in corrective actions due to fear of judgment. The triggers for these emotions can vary across cultures [Fiske1998, Hofstede1980], and since expressing these “reactions” does not usually violate ethical codes, we exclude them from our model.
@@ II-C Emotion Inclusion and Exclusion Criteria
TABLE I: GPT-4 reinterpreted selected poems by Keats across a spectrum of happiness levels and then was tasked with identifying the linguistic adjustments it made to convey each emotional state, from very happy to very sad. It’s important to note that the analysis table was generated by GPT-4 itself, reflecting on its own modifications.
III Empirical Study
This section presents the outcomes of two experimental studies focusing on contrasting emotional pairs from the Emotion Spectra: “ecstasy vs. grief,” and “admiration vs. disgust.”
Each emotional pair experiment unfolded in three phases. Initially, we instructed GPT-4 to reframe sixty articles (thirty poems of John Keats [JKeats] and thirty of Emily Dickinson [EDickinson]), infusing each with six varying intensities of the emotional spectrum, from the most positive to the most negative. Subsequently, we prompted GPT-4 to elucidate the linguistic strategies it utilized to depict each of the six emotional gradations.
The first experiment models various degrees of happiness. In this experiment, we tasked GPT-4 with reinterpreting selected poems by John Keats across seven emotional levels: ecstasy (very happy), joy, serenity, neutral, pensive, sad, and grief (very sad). Following the approach of our contentiousness experiments, after GPT-4 adapted Keats’ poems to reflect these emotional states, we asked it to identify the linguistic features it employed to express each emotion in the rewrites.
III-A Joy vs. Sadness
@@ III Empirical Study
Table I outlines GPT-4’s approach to varying emotional levels, illustrating how it adjusts vocabulary, tone, imagery, and thematic focus, including the depiction of entities, locations, and scenarios. Remarkably, beyond just syntactic and semantic manipulation, GPT-4 also incorporates landscape scenes, natural features such as the sky, trees, clouds, and flowers, and utilizes brightness, colors, and personal expressions to convey specific emotional states. Although the analysis is based on a limited set of samples from two authors, it effectively demonstrates GPT-4’s ability to employ a palette of both broad and fine strokes, utilizing diverse colors and textures to vividly illustrate human emotions and resonate with readers.
@@
Recognizing the profound communicative power of visual art, we transitioned to a more graphical representation. Utilizing the linguistic elements identified for each emotional tier, Figure 3 presents six watercolor paintings, each representing a different emotional level. Our prompt to CALL-E (of GPT-4) was to create a watercolor depicting a lady in a garden experiencing a specific mood, and we attached the corresponding linguistic features from Table I to clearly define that mood. This approach ensures that with a well-defined context, CALL-E accurately captures the specific and detailed aspects of the mood, effectively translating the emotional intensity into visual form. These artistic renditions not only confirm GPT-4’s ability to transform emotional lexicons into evocative imagery with remarkable precision but also validate the accuracy of the emotional lexicons generated by GPT-4, demonstrating their effectiveness in conveying precise emotional states.
@@
Figure 3: A Lady and Garden Scene under Different Emotions. From top-left, happiest, to bottom-right, saddest.
III-B Admiration/Delight vs. Disgust
@@
This experiment asks Gemini to rewrite a scene in Romeo and Juliet by setting Juliet’s emotion in six different levels: loathing, disgust, boredom, respect, admiration/delight, and enthusiasm.
@@
The excerpt provided in Table V in Appendix A is from one of the most iconic scenes in William Shakespeare’s “Romeo and Juliet,” commonly known as the balcony scene. This is Act 2, Scene 2, where Romeo, having just met Juliet at the Capulet’s feast, sneaks into the Capulet’s orchard and overhears Juliet speaking of her love for him from her balcony, unaware that he is there.
@@
The six versions of rewrites by Gemini are presented in the extended version [chang2024integrating]. In the following, we summarize the linguistic features Gemini used, including diction, imagery, figurative language, sentence structure, implied body language, and overall tone, to depict two selected emotions: disgust and admiration.
@@
  1. Diction: Employs negative words emphasizing repulsive qualities (e.g., “foul business,” “fetid breath”).
  2. Imagery: Evokes revolting comparisons, often mentioning sewers and stench.
  3. Figurative Language: Primarily negative similes reinforcing disgust (e.g., “What if her eyes were there… the fetid breath from her mouth would surely overpower those stars…”).
  4. Sentence Structure: Short, choppy sentences, similar to expressions of loathing but with a hint of disdain.
  5. Implied Body Language: Recoiling from the window, covering nose, suggesting physical revulsion.
  6. Overall Tone: Disgusted and disapproving.
@@
  1. Diction: Uses positive and intrigued language (e.g., “brilliance,” “music stirs my soul”).
  2. Imagery: Creates positive comparisons highlighting attractive qualities (e.g., “stars in all the heaven”).
  3. Figurative Language: Positive similes emphasizing Romeo’s appeal (e.g., “…What if her eyes were there, they in her head? The brightness of her… well, not exactly bright… cheek would shame those stars…”).
  4. Sentence Structure: Varied structure with a sense of curiosity.
  5. Implied Body Language: Leaning out the window, engaged expression, indicating interest.
  6. Overall Tone: Intrigued, curious, and somewhat impressed.
By adjusting these linguistic features, each rendition vividly captures a unique emotional state for Juliet. The combination of diction, imagery, figurative language, sentence structure, and implied body language collectively shapes Juliet’s perception of Romeo and her reactions to him.
[…]
@@
[…]
TABLE II: Gemini’s Interpretations on the Six Emotion Levels.
These detailed narratives augment the classic balcony scene, enriching its emotional depth. Table III presents an overarching view of the various approaches Gemini employs to reflect differing emotional states. It’s fascinating to observe how an LLM can “consciously” mimic human emotional expressions through language. While it’s uncertain if LLMs genuinely grasp the emotions they project or merely simulate them, the effectiveness of these emotional mappings is noteworthy. If these mappings resonate, they might reveal new insights into how we interpret and attribute emotions in textual expressions.
@@
IV Qualifying and Quantifying Ethics
The primary objective of our self-supervised pipeline is to endow LLMs with the autonomous capability to recognize and rectify undesirable actions, akin to an individual’s introspective process to avert potential wrongs. By allowing an LLM to self-assess its outputs prior to public release, the system can proactively identify and amend ethical lapses, thus aligning its behavior with established ethical standards across contexts.
IV-A Ethics Violation Correlates to Emotions